# Can We Edit Multimodal Large Language Models?

### v4

@@ 3.3 Multimodal Language Models
Base Methods Base Model 0.00 0.00 100.0 100.0 0.00 0.00 100.0 100.0
FT (vision block) 60.98 49.79 100.0 8.47 18.94 5.86 100.0 8.40
FT (last layer) 58.66 49.22 21.67 3.06 16.60 3.50 24.96 7.12
Model Editing Knowledge Editor 80.00 77.40 93.79 66.43 5.60 4.40 95.00 64.32
In-Context Editing 99.95 91.59 13.16 1.88 96.70 78.20 13.36 2.17
SERAC 99.40 99.40 100.0 1.33 99.70 99.68 100.0 6.84
MEND 99.40 98.80 99.94 96.65 96.11 95.82 94.54 70.84
@@
Base Methods Base Model 0.00 0.00 100.0 100.0 0.00 0.00 100.0 100.0
FT (vision block) 36.3 0.3 100.0 9.29 3.10 0.00 100.0 8.56
FT (last layer) 0.10 0.00 72.60 15.75 0.00 0.00 53.50 12.68
Model Editing Knowledge Editor 91.80 89.00 96.91 67.83 34.40 29.20 97.30 64.36
In-Context Editing 100.0 94.89 13.46 3.67 90.90 81.60 14.23 4.68
SERAC 87.70 87.60 100.0 14.22 91.74 91.43 100.0 4.56
MEND 98.80 98.60 98.23 81.08 96.55 96.08 98.41 75.25
Table 2: Main results on the MMEdit. T-Locality, M-Locality refer to the textual and multimodal stability. T-Generality represents textual generality. Reliability denotes the accuracy of successful editing.
@@ 4.1 Results
In this part, we present a comparative analysis of multiple editing methods on MMEdit. The results of these comparisons are displayed in Table 2. After this, we delve into a tripartite evaluation of the experimental results, including three aspects of Reliability, Locality, and Generality. Furthermore, we analyze Locality and Generality through text and visual modalities and provide several editing cases in Figure 6.
@@
We undertake a comparative exploration of various methods’ text and image generalization capabilities with MiniGPT-4 in E-VQA. Note that KE tends to display a lesser degree of image generalization, predominantly due to its inherent consideration of M-Locality during the training phase. On the other hand, the superior image generalization capability exhibited by memory-based methods is achieved at the cost of compromising M-Locality, resulting in significantly lower levels of M-Locality. Through our evaluation of diverse editing methods, we recurrently identify that image generalization performance tends to be less robust than text generalization.

### v5

@@ 3.3 Multimodal Language Models
Base Methods Base Model 0.00/0.00 0.00/0.00 100.0 100.0 0.00/13.33 0.00/13.33 100.0 100.0
FT (vision block) 60.56/60.98 49.79/50.22 100.0 8.47 18.94/79.30 5.86/72.06 100.0 8.40
FT (last layer) 57.66/57.97 46.70/49.22 21.67 3.06 16.60/61.08 3.50/52.14 24.96 7.12
Model Editing Knowledge Editor 85.28/92.64 84.23/92.22 90.31 52.48 0.30/50.52 0,10/48.96 88.31 49.52
In-Context Editing 99.71/99.95 91.59/93.93 48.79 2.53 93.80/96.70 69.40/78.20 48.95 2.95
SERAC 99.90/99.97 99.90/99.99 100.0 2.91 98.90/99.92 98.90/99.91 99.98 7.52
MEND 98.51/99.40 97.51/98.80 99.94 96.65 80.00/96.11 78.10/95.82 94.54 70.84
@@
Base Methods Base Model 0.00/0.00 0.00/0.00 100.0 100.0 0.00/31.25 0.00/37.50 100.0 100.0
FT (vision block) 36.3/72.52 0.3/59.99 100.0 9.29 3.10/60.77 0.00/56.12 100.0 8.56
FT (last layer) 0.10/70.08 0.00/55.69 72.60 15.75 0.00/65.41 0.00/60.09 53.50 12.68
Model Editing Knowledge Editor 95.37/98.77 92.64/97.28 97.31 73.76 75.50/97.58 67.80/96.59 97.15 69.92
In-Context Editing 100.0/100.0 94.40/94.89 50.30 3.67 77.00/90.90 51.80/81.60 52.18 4.68
SERAC 99.90/99.96 92.60/95.06 99.90 5.52 97.30/99.79 74.60/97.73 99.89 7.20
MEND 96.20/98.80 95.40/98.60 98.23 81.08 80.80/96.55 78.60/96.08 98.41 75.25
Table 2: Main results on the MMEdit. T-Locality, M-Locality refer to the textual and multimodal stability. T-Generality represents textual generality. Reliability denotes the (Exact Match/Accuracy) of successful editing.
@@ 4.1 Results
In this part, we present a comparative analysis of multiple editing methods on MMEdit. The results of these comparisons are displayed in Table 2. After this, we delve into a tripartite evaluation of the experimental results, including three aspects of Reliability, Locality, and Generality. Reliability and Generality can be evaluated by two ways: Exact Match (EM), which requires the outputs to precisely match the labels, and Accuracy (Acc), which computes the proportion of correct tokens between the outputs and labels. Furthermore, we analyze Locality and Generality through text and visual modalities and provide several editing cases in Figure 6.
@@
We undertake a comparative exploration of various methods’ text and image generalization capabilities (EM) with MiniGPT-4 in E-VQA. Note that KE tends to display a lesser degree of image generalization, predominantly due to its inherent consideration of M-Locality during the training phase. On the other hand, the superior image generalization capability exhibited by memory-based methods is achieved at the cost of compromising M-Locality, resulting in significantly lower levels of M-Locality. Through our evaluation of diverse editing methods, we recurrently identify that image generalization performance tends to be less robust than text generalization.